Evaluate lim x tends to 1
root 1-cos2(x-1)/(x-1)

Grade:11

3 Answers

orange panda
11 years ago

x-1=h(let). so lim h tends to zero sinh(2^0.5)/h=root 2

chandradeep pokhariya
11 years ago

Limx-1 √(1-cos2(x-1))/x-1

Let x-1->0 so taking x-1=h

Also converting cos2(h) in 1-2sin2

√2 is the answer

Gautham Pai M K
11 years ago

 Good Luck!!